OpenStudy (anonymous):

Monica purchased adult and youth tickets for the fair. She bought x adult tickets for $28 each and y youth tickets for $15. Write an expression that can be used to show the total cost, C, of all tickets.

OpenStudy (anonymous):

The options for the question are: C=(28+15)x C=(28+15)(x+y) C=28x+15y C=28y+15x

OpenStudy (anonymous):

read the problem critically. How many are associated with x? How many are linked to y?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

I am waiting for you to answer the question: What number is by the letter x?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

yes, so place the x after the 28, and tell me what else it says

OpenStudy (anonymous):

well the y is next to the 15

OpenStudy (anonymous):

that is correct. So how do we show it?

OpenStudy (anonymous):

I'm thinking the right answer is either b or c its just they are both different properties

OpenStudy (anonymous):

Let's take a moment. Where did I ask you to place the x? In front or behind the number?

OpenStudy (anonymous):

in front

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Reread what I wrote

OpenStudy (anonymous):

it says after

OpenStudy (anonymous):

isn't that in front?

OpenStudy (anonymous):

yes, so write the x after the number

OpenStudy (anonymous):

so its c right?

OpenStudy (anonymous):

yes, C is correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are welcome